# Properties of kaon-condensed phase in hyperon-mixed matter with three-baryon forces

### v1

@@ II.2 Estimation of the $K$-$B$ sigma terms
Recent lattice QCD results suggest small $\bar{s}s$ condensate in the nucleon, i. e., $y_{N}\equiv 2\langle N|\bar{s}s|N\rangle/\langle N|(\bar{u}u+\bar{d}d)|N\rangle$ = 0.03 -- 0.2 ohki08 ; Durr2016 ; Alex2020 . In particular, for $\sigma_{s}$ = 0, one can see from Eq. (22) that $\sigma_{0}=\Sigma_{\pi N}\simeq$ 25 MeV. This value of $\Sigma_{\pi N}$ is too small compared to the phenomenological values (40--60) MeV gls1991 ; a2021 which are deduced from the analyses of $\pi$-$N$ scattering and pionic atoms, or lattice QCD results $\sim$ 40 MeV a2021 . It has been shown that nonlinear effects beyond chiral perturbation can make both the value of $\Sigma_{\pi N}$ and the octet baryon mass splittings consistent with experiments with a small strangeness content of the proton a2021 ; jk1987 ; hk1991 . Here we take into account the nonlinear effect on the quark condensates which originate from the additional rest mass contribution of baryons, $\Delta M(m_{s})$ in higher order with respect to $m_{s}$ (see the Appendix X.1). Then the $Kb$ sigma terms (19) are modified by the replacement: $a_{3}\rightarrow\widetilde{a}_{3}\equiv a_{3}-\Delta/4$ with $\Delta\equiv\partial\Delta M/\partial m_{s}$. The nonlinear effect $\Delta$ is absorbed into $\widetilde{a}_{3}$.
To estimate the allowable value of the $\Sigma_{KN}$, we start with the general relation among $\Sigma_{KN}$, $\Sigma_{\pi N}$, and $\sigma_{s}$, or $y_{N}$,
@@
The standard value for $\Sigma_{\pi N}$ has been taken to be $\sim$ 45 MeV phenomenologically. Recently the higher values (50--60) MeV are suggested from the phenomenological analyses a2021 . In view of this, reading off from Fig. 1, we take two cases of $\Sigma_{Kn}$ = 300 MeV with $y_{N}$ = 0 and 400 MeV with $y_{N}$ = 0.2 as typical values for $\Sigma_{Kn}$ throughout this paper. The corresponding quantities, $\Sigma_{\pi N}$, $\Delta$, and $\Sigma_{Kb}$ ($b=p,\Lambda,\Sigma^{-},\Xi^{-}$) together with $a_{3}$, $\widetilde{a}_{3}$ are also determined for fixed values of ($m_{u}$, $m_{d}$, $m_{s}$) and ($a_{1}$, $a_{2}$). The result is listed in Table 1.
@@
Recent lattice simulation results on the current quark masses give rather smaller values, $(m_{u},m_{d},m_{s})$ = (2.2, 4.7, 95) MeV PDG2020 , than adopted in this paper. However, the ratio $(m_{u}+m_{s})/2\hat{m}$ ($\sim$14) is almost the same in both set of quark masses. Therefore, one can see from (25) that the $\Sigma_{KN}$ is little altered by the use of different quark masses once the $\Sigma_{\pi N}$ is given. It can also be shown that the least-square fitting for the empirical octet mass splitting by the use of the lighter quark masses little changes the result of not only $\Sigma_{KN}$ but also $\Sigma_{KY}$, as seen from Table 1.
It is to be noted that the pion-baryon sigma terms [$\sigma_{bq}=\hat{m}\langle b|(\bar{u}u+\bar{d}d)|b\rangle$] and strangeness sigma terms [$\sigma_{bs}=m_{s}\langle b|\bar{s}s|b\rangle$] in the octet baryons ($b$) have been derived from analyses of the lattice QCD simulations for the octet baryon masses sty2013 ; Lutz2014 . The relative ordering of $\sigma_{bq}$, $\sigma_{bs}$ for $b=(N,Y(=\Lambda,\Sigma^{-},\Xi^{-}))$ estimated in our model agrees well with those in Ref. sty2013 , while there is a little difference in absolute values for $\sigma_{Yq}$, $\sigma_{Ys}$ between our results and those in Fig. 2 in sty2013 .
@@ III Baryon interactions
In our framework, the baryon interactions are separated into two parts: (I) Two-body $B$-$B$ interaction which may be relevant mainly to the intermediate and long-range regions ($\rho_{\rm B}\lesssim 2\rho_{0}$), where baryons can be considered to be point-like and the $B$-$B$ interaction is mainly mediated by the scalar ($\sigma$, $\sigma^{\ast}$) mesons and vector ($\omega$, $\rho$, $\phi$) mesons in the RMF. (II) Phenomenological three-body forces. One is the three-nucleon repulsion (TNR) and TNA, both of which are important to reproduce the saturation of the SNM around $\rho_{0}$. Another is the three-baryon repulsion (TBR), which is an extension of the TNR to high-densities where hyperons as well as nucleons are occupied in matter. TBR is relevant to the short-range region, where quark structure of baryon reveals itself, and the origin of short-range repulsion is connected with the quark confinement mechanism which is spin-flavor independent.
@@ III.2 Three-baryon repulsive force
It may be naturally understood that the three-body repulsion is qualitatively independent on spin-flavor of baryons, reflecting the confinement mechanisms of quarks at high-density region. Thus it is assumed to work universally between any baryon species. Along with this viewpoint, we adopt a specific model for the UTBR proposed by Tamagaki based on the string-junction model (SJM)  t2008 ; tnt2008 . We utilize the density-dependent effective two-body potential $U_{\rm SJM}(1,2;\rho_{\rm B})$ between baryons 1 and 2, by integrating out variables of the third baryon participating the UTBR:
@@
with $W_{0}$ being the strength of the order of $B$-antibaryon ($\bar{B}$) excitation energy ($\simeq$ 2 GeV) t2008 , and $g({\bf r}_{i}-{\bf r}_{j})$ being the wavefunction between $B_{i}$ and $B_{j}$. Taking the wavefunction $g({\bf r})$ as the Gaussian form, $g({\bf r})=\exp[-(r/\eta_{c})^{2}]$ with $r=|{\bf r}_{i}-{\bf r}_{j}|$ and $\eta_{c}$ [= (0.45$\sim$ 0.50) fm] being the range of the repulsive core for baryon forces, one obtains
@@
with $j_{0}(qr)=\sin(qr)/(qr)$ being the spherical Bessel function of the 1st kind. The $f_{\rm src}(r)$ is used as the statistical weighted average of the ${}^{1}{\rm E}$ and ${}^{3}{\rm O}$ wave functions as $f_{\rm src}(r)\equiv[f_{\rm src}(^{1}{\rm E})(r)+3f_{\rm src}(^{3}{\rm O})(r)]/4$, where we, for simplicity, substitute for $f_{\rm src}(r)$ the result of the reaction matrix calculation in neutron matter by the use of the one-pion exchange Gaussian (OPEG)-A potential assuming that the density-dependence is weak t2008 ; tnt2008 :
@@ III.3 Three-nucleon attractive force
To simulate the attractive contribution from the TNA to the binding energy for $\rho_{\rm B}\lesssim\rho_{0}$ , we adopt the density-dependent effective two-body potential by Nishizaki, Takatsuka and Hiura (NTH 1994) nth1994 , which was phenomenologically introduced and the direct term of which agrees with the expression by Lagaris and Pandharipande (LP1981) lp1981 :
@@ IV Description of the ground state for the ($Y$+$K$) phase
Our interaction model in the MRMF consists of the Lagrangian density ${\cal L}_{K}$ +${\cal L}_{B,M}$ [Eqs. (14) and (35)]. In addition, the three-baryon forces UTBR and TNA are taken into account as the effective two-baryon potentials [Eqs. (43) and (44)].
@@ IV.1 Energy density expression for the ($Y$+$K$) phase
$=$ $\frac{\pi^{3/2}}{2}\rho_{\rm B}U_{\rm SJM2}(r=0;\rho_{\rm B})\cdot(\widetilde{\lambda}_{r})^{3}\ ,$ (52)
@@
$C_{b}(r)\equiv\int_{|{\bf k}_{b}|\leq k_{F}(b)}\frac{d^{3}|{\bf k}_{b}|}{(2\pi)^{3}}e^{-i{\bf k}_{b}\cdot{\bf r}}=\frac{3}{2}\rho_{b}\frac{j_{1}(k_{F}(b)r)}{k_{F}(b)r}$ (54)
with $j_{1}(x)$ [$\equiv(\sin x-x\cos x)/x^{2}$] being the spherical Bessel function of the 1st kind, $\rho_{b}$ and $k_{F}(b)$ the number density and the Fermi momentum of baryon $b$ ($b=p,n,\Lambda,\Sigma^{-},\Xi^{-}$), respectively. The exchange contribution (53) to the energy density is shown to be numerically small in comparison with the direct term (52) except for $r\sim 0$. Therefore, it is reasonable to neglect the exchange contribution to the energy density in line with the Hartree approximation in the RMF framework.
@@
with $\gamma_{a}\equiv(\pi^{3/2}/2)V_{a}\lambda_{a}^{3}$ and $x_{p}=\rho_{p}/\rho_{\rm B}$ the proton-mixing ratio. The exchange term of the TNA has been shown to be very small in comparison with (57nth1994 , so we also neglect it.
@@ V Properties of symmetric nuclear matter and pure neutron matter
Here we elucidate the ground state properties of SNM in the (MRMF+UTBR+TNA) model and discuss the effects of the three-nucleon forces, UTBR and TNA, on the stiffness of the EOS of SNM at densities around and beyond $\rho_{0}$. The EOS of pure neutron matter (PNM) is also briefly mentioned.
@@ V.1 Meson-nucleon coupling constants from saturation properties in SNM
With Eqs. (49)--(57), the total energy per nucleon, $E$(total), in SNM is written as $E$(total)=$E$(two-body)+$E$(TNR)+$E$(TNA) = [${\cal E}$(${\cal E}$(${\cal E}$(TNA)]/$\rho_{\rm B}$, where
${\cal E}$ $({\rm two-}{\rm body})={\cal E}_{B,M}/\rho_{\rm B}$ (81)
$=$ $\sum_{N=p,n}\frac{2}{(2\pi)^{3}}\int_{|{\bf p}|\leq p_{F}}d^{3}|{\bf p}|(|{\bf p}|^{2}+M_{N}^{\ast 2})^{1/2}$ (82)
@@
$m_{\sigma}^{2}\langle\sigma\rangle_{0}$ $=$ $g_{\sigma N}\rho_{N}^{s}|_{\rho_{\rm B}=\rho_{0}}$ (86)
@@
where $\rho_{N}^{s}$ (=$\rho_{p}^{s}+\rho_{n}^{s}$) is the nuclear scalar density. The meson-nucleon coupling constants, $g_{\sigma N}$, $g_{\omega N}$, $g_{\rho N}$, the meson mean-fields at $\rho_{0}$, $\langle\sigma\rangle_{0}$, $\langle\omega_{0}\rangle_{0}$, and the parameters $\gamma_{a}$, $\eta_{a}$ in TNA associated with isospin symmetry are obtained from Eqs. (85) and (87), together with the empirical values of the incompressibility $K_{0}$=240 MeV GC2018 , the symmetry energy $S_{0}$ (=31.5 MeV) at $\rho_{0}$ lh2013 , and the slope $L$ of the symmetry energy, being defined as $L$$\equiv 3\rho_{0}\left(dS(\rho_{\rm B})/d\rho_{B}\right)_{\rho_{B}=\rho_{0}}=3P_{\rm PNM}(\rho_{0})/\rho_{0}$, where $S(\rho_{\rm B})$ is the symmetry energy and $P_{\rm PNM}(\rho_{\rm B})$ is the pressure for pure neutron matter (PNM); both are given as functions of baryon density. The expressions for $K_{0}$, $S_{0}$, and $L$ in the present model are given by Eqs. (121), (124), and (127), respectively in Appendix X.2. Allowable values of $S_{0}$ and $L$ have been constrained from analyses of terrestrial experiments such as measurements of giant dipole resonances, dipole polarizabilities, neutron skins of neutron-rich nuclei, and Heavy-Ion Collisions, etc. Up to now, the empirical value of the $L$ has a large uncertainty, ranging from 30 MeV to 90 MeV lp2016 ; oertel2017 ; xia2021 . Recent measurements of neutron skin thickness from scattering of the spin-polarized electrons on 208Pb by the PREX-II experiments adhikari2021 and on 48C by the CREX adhikari2022 lead to controversy about the value of $L$ reed2021 ; Reinhard2021 ; reed2024 . Another experiment is the measurement of the spectra of charged pions produced by colliding rare isotope Sn beams with isotopically enriched Sn targets. The deduced value of $L$ has been shown to be 42 $<L<$ 117 MeV estee2021 . Also there are attempts to get information on $L$ from the astronomical observations associated with $X$-ray bursters sotani2018 . In the present work, the parameters, $V_{r}$, $c_{r}$, and $\lambda_{r}$ in $U_{\rm SJM2}$ [(43)] are fixed to be those used in SJM2 model in Ref. tnt2008 , so that the $E$ (TNR) is commonly given irrespective of choice of the other parameters. Allowable values of the $L$ are chosen so that the overall density-dependence of the energy contributions, $E$ (two-body), $E$ (TNA), and $E$ (total) around $\rho_{0}$ in SNM are similar to those obtained by the standard variational calculation by the use of the $v_{14}$ two-body potential with addition of the phenomenological TNR and TNA by Lagaris-Pandahripande [LP (1981)] lp1981 . It turns out that $E$ (two-body) in our model is not bound for $L\lesssim 55$ MeV and that its density-dependence is quite different from that of LP (1981). For $L\gtrsim 75$ MeV, an onset density of KC exceeds a central density of neutron star with maximum mass. Hence we take the typical three values $L$ = (60, 65, 70) MeV throughout this paper. In Table 2, the relevant quantities associated with the (MRMF+UTBR+TNA) model are listed for three cases of $L$ = (60, 65, 70) MeV.

### v2

@@ II.2 Estimation of the $K$-$B$ sigma terms
Recent lattice QCD results suggest small $\bar{s}s$ condensate in the nucleon, i. e., $y_{N}\equiv 2\langle N|\bar{s}s|N\rangle/\langle N|(\bar{u}u+\bar{d}d)|N\rangle$ = 0.03 -- 0.2 ohki08 ; Durr2016 ; Alex2020 . In particular, for $\sigma_{s}$ = 0, one can see from Eq. (22) that $\sigma_{0}=\Sigma_{\pi N}\simeq$ 25 MeV. This value of $\Sigma_{\pi N}$ is too small compared to the phenomenological values (40--60) MeV gls1991 ; a2021 which are deduced from the analyses of $\pi$-$N$ scattering and pionic atoms, or lattice QCD results $\approx$ 40 MeV a2021 . It has been shown that nonlinear effects beyond chiral perturbation can make both the value of $\Sigma_{\pi N}$ and the octet baryon mass splittings consistent with experiments with a small strangeness content of the proton a2021 ; jk1987 ; hk1991 . Here one takes into account the nonlinear effect on the quark condensates which originate from the additional rest mass contribution of baryons, $\Delta M(m_{s})$ in higher order with respect to $m_{s}$ (see the Appendix X.1). Then the $Kb$ sigma terms (19) are modified by the replacement: $a_{3}\rightarrow\widetilde{a}_{3}\equiv a_{3}-\Delta/4$ with $\Delta\equiv\partial\Delta M/\partial m_{s}$. The nonlinear effect $\Delta$ is absorbed into $\widetilde{a}_{3}$.
To estimate the allowable value of the $\Sigma_{KN}$, one starts with the general relation among $\Sigma_{KN}$, $\Sigma_{\pi N}$, and $\sigma_{s}$, or $y_{N}$,
@@
The standard value for $\Sigma_{\pi N}$ has been taken to be $\approx$ 45 MeV phenomenologically. Recently the higher values (50--60) MeV are suggested from the phenomenological analyses a2021 . In view of this, reading off from Fig. 1, one takes two cases of $\Sigma_{Kn}$ = 300 MeV with $y_{N}$ = 0 and 400 MeV with $y_{N}$ = 0.2 as typical values for $\Sigma_{Kn}$ throughout this paper. The corresponding quantities, $\Sigma_{\pi N}$, $\Delta$, and $\Sigma_{Kb}$ ($b=p,\Lambda,\Sigma^{-},\Xi^{-}$) together with $a_{3}$, $\widetilde{a}_{3}$ are also determined for fixed values of ($m_{u}$, $m_{d}$, $m_{s}$) and ($a_{1}$, $a_{2}$). The result is listed in Table 1.
@@
Recent lattice simulation results on the current quark masses give rather smaller values, $(m_{u},m_{d},m_{s})$ = (2.2, 4.7, 95) MeV PDG2020 , than adopted in this paper. However, the ratio $(m_{u}+m_{s})/2\hat{m}$ ($\approx$14) is almost the same in both set of quark masses. Therefore, one can see from (25) that the $\Sigma_{KN}$ is little altered by the use of different quark masses once the $\Sigma_{\pi N}$ is given. It can also be shown that the least-square fitting for the empirical octet mass splitting by the use of the lighter quark masses little changes the result of not only $\Sigma_{KN}$ but also $\Sigma_{KY}$, as seen from Table 1.
It is to be noted that the pion-baryon sigma terms [$\sigma_{bq}=\hat{m}\langle b|(\bar{u}u+\bar{d}d)|b\rangle$] and strangeness sigma terms [$\sigma_{bs}=m_{s}\langle b|\bar{s}s|b\rangle$] in the octet baryons ($b$) have been derived from analyses of the lattice QCD simulations for the octet baryon masses sty2013 ; Lutz2014 . The relative ordering of $\sigma_{bq}$, $\sigma_{bs}$ for $b=(N,Y(=\Lambda,\Sigma^{-},\Xi^{-}))$ estimated in the present model agrees well with those in Ref. sty2013 , while there is a little difference in absolute values for $\sigma_{Yq}$, $\sigma_{Ys}$ between the present results and those in Fig. 2 in sty2013 .
@@ III Baryon interactions
In the framework in this paper, the baryon interactions are separated into two parts: (I) Two-body $B$-$B$ interaction which may be relevant mainly to the intermediate and long-range regions ($\rho_{\rm B}\lesssim 2\rho_{0}$), where baryons can be considered to be point-like and the $B$-$B$ interaction is mainly mediated by the scalar ($\sigma$, $\sigma^{\ast}$) mesons and vector ($\omega$, $\rho$, $\phi$) mesons in the RMF. (II) Phenomenological three-body forces. One is the three-nucleon repulsion (TNR) and TNA, both of which are important to reproduce the saturation of the SNM around $\rho_{0}$. Another is the three-baryon repulsion (TBR), which is an extension of the TNR to high-densities where hyperons as well as nucleons are occupied in matter. TBR is relevant to the short-range region, where quark structure of baryon reveals itself, and the origin of short-range repulsion is connected with the quark confinement mechanism which is spin-flavor independent.
@@ III.2 Three-baryon repulsive force
It may be naturally understood that the three-body repulsion is qualitatively independent on spin-flavor of baryons, reflecting the confinement mechanisms of quarks at high-density region. Thus it is assumed to work universally between any baryon species. Along with this viewpoint, one adopts a specific model for the UTBR proposed by Tamagaki based on the string-junction model (SJM)  t2008 ; tnt2008 . One utilizes the density-dependent effective two-body potential $U_{\rm SJM}(1,2;\rho_{\rm B})$ between baryons 1 and 2, by integrating out variables of the third baryon participating the UTBR:
@@
with $W_{0}$ being the strength of the order of $B$-antibaryon ($\bar{B}$) excitation energy ($\simeq$ 2 GeV) t2008 , and $g({\bf r}_{i}-{\bf r}_{j})$ being the wavefunction between $B_{i}$ and $B_{j}$. Taking the wavefunction $g({\bf r})$ as the Gaussian form, $g({\bf r})=\exp[-(r/\eta_{c})^{2}]$ with $r=|{\bf r}_{i}-{\bf r}_{j}|$ and $\eta_{c}$ [= (0.45 -- 0.50) fm] being the range of the repulsive core for baryon forces, one obtains
@@
with $j_{0}(qr)=\sin(qr)/(qr)$ being the spherical Bessel function of the 1st kind. The $f_{\rm src}(r)$ is used as the statistical weighted average of the ${}^{1}{\rm E}$ and ${}^{3}{\rm O}$ wave functions as $f_{\rm src}(r)\equiv[f_{\rm src}(^{1}{\rm E})(r)+3f_{\rm src}(^{3}{\rm O})(r)]/4$, where, for simplicity, the result of the reaction matrix calculation in neutron matter by the use of the one-pion exchange Gaussian (OPEG)-A potential is substituted for $f_{\rm src}(r)$, by the assumption that the density-dependence is weak t2008 ; tnt2008 :
@@ III.3 Three-nucleon attractive force
To simulate the attractive contribution from the TNA to the binding energy for $\rho_{\rm B}\lesssim\rho_{0}$ , one adopts the density-dependent effective two-body potential by Nishizaki, Takatsuka and Hiura (NTH 1994) nth1994 , which was phenomenologically introduced and the direct term of which agrees with the expression by Lagaris and Pandharipande (LP1981) lp1981 :
@@ IV Description of the ground state for the ($Y$+$K$) phase
The present interaction model in the MRMF consists of the Lagrangian density ${\cal L}_{K}$ +${\cal L}_{B,M}$ [Eqs. (14) and (35)]. In addition, the three-baryon forces UTBR and TNA are taken into account as the effective two-baryon potentials [Eqs. (43) and (44)].
@@ IV.1 Energy density expression for the ($Y$+$K$) phase
$=$ $\frac{\pi^{3/2}}{2}\rho_{\rm B}^{2}U_{\rm SJM2}(r=0;\rho_{\rm B})\cdot(\widetilde{\lambda}_{r})^{3}\ ,$ (52)
@@
$C_{b}(r)\equiv\int_{|{\bf k}_{b}|\leq p_{F}(b)}\frac{d^{3}|{\bf k}_{b}|}{(2\pi)^{3}}e^{-i{\bf k}_{b}\cdot{\bf r}}=\frac{3}{2}\rho_{b}\frac{j_{1}(p_{F}(b)r)}{p_{F}(b)r}$ (54)
with $j_{1}(x)$ [$\equiv(\sin x-x\cos x)/x^{2}$] being the spherical Bessel function of the 1st kind, $\rho_{b}$ and $p_{F}(b)$ the number density and the Fermi momentum of baryon $b$ ($b=p,n,\Lambda,\Sigma^{-},\Xi^{-}$), respectively. The exchange contribution (53) to the energy density is shown to be numerically small in comparison with the direct term (52) except for $r\approx 0$. Therefore, it is reasonable to neglect the exchange contribution to the energy density in line with the Hartree approximation in the RMF framework.
@@
with $\gamma_{a}\equiv(\pi^{3/2}/2)V_{a}\lambda_{a}^{3}$ and $x_{p}=\rho_{p}/\rho_{\rm B}$ the proton-mixing ratio. The exchange term of the TNA has been shown to be very small in comparison with (57nth1994 , so it is also neglected.
@@ V Properties of symmetric nuclear matter and pure neutron matter
Here the ground state properties of SNM in the (MRMF+UTBR+TNA) model are elucidated, and the effects of the three-nucleon forces, UTBR and TNA, on the stiffness of the EOS of SNM at densities around and beyond $\rho_{0}$ are discussed. The EOS of pure neutron matter (PNM) is also briefly mentioned.
@@ V.1 Meson-nucleon coupling constants from saturation properties in SNM
With Eqs. (49)--(57), the total energy per nucleon, $E$(total) in SNM is written as $E$(total)=$E$(two-body)+$E$(TNR)+$E$(TNA) = [${\cal E}$(${\cal E}$(${\cal E}$(TNA)]/$\rho_{\rm B}$, where
${\cal E}({\rm two-}{\rm body})$ $=$ $\sum_{N=p,n}\frac{2}{(2\pi)^{3}}\int_{|{\bf p}|\leq p_{F}}d^{3}|{\bf p}|(|{\bf p}|^{2}+M_{N}^{\ast 2})^{1/2}$ (81)
@@
$m_{\sigma}^{2}\langle\sigma\rangle_{0}$ $=$ $g_{\sigma N}\rho_{N}^{s}|_{\rho_{\rm B}=\rho_{0}}\ ,$ (85)
@@
where $\rho_{N}^{s}$ (=$\rho_{p}^{s}+\rho_{n}^{s}$) is the nuclear scalar density. The meson-nucleon coupling constants, $g_{\sigma N}$, $g_{\omega N}$, $g_{\rho N}$, the meson mean-fields at $\rho_{0}$, $\langle\sigma\rangle_{0}$, $\langle\omega_{0}\rangle_{0}$, and the parameters $\gamma_{a}$, $\eta_{a}$ in TNA associated with isospin symmetry are obtained from Eqs. (84) and (86), together with the empirical values of the incompressibility $K_{0}$=240 MeV GC2018 , the symmetry energy $S_{0}$ (=31.5 MeV) at $\rho_{0}$ lh2013 , and the slope $L$ of the symmetry energy, being defined as $L$$\equiv 3\rho_{0}\left(dS(\rho_{\rm B})/d\rho_{B}\right)_{\rho_{B}=\rho_{0}}=3P_{\rm PNM}(\rho_{0})/\rho_{0}$, where $S(\rho_{\rm B})$ is the symmetry energy and $P_{\rm PNM}(\rho_{\rm B})$ is the pressure for pure neutron matter (PNM); both are given as functions of baryon density. The expressions for $K_{0}$, $S_{0}$, and $L$ in the present model are given by Eqs. (120), (123), and (126), respectively in Appendix X.2. Allowable values of $S_{0}$ and $L$ have been constrained from analyses of terrestrial experiments such as measurements of giant dipole resonances, dipole polarizabilities, neutron skins of neutron-rich nuclei, and Heavy-Ion Collisions, etc. Up to now, the empirical value of the $L$ has a large uncertainty, ranging from 30 MeV to 90 MeV lp2016 ; oertel2017 ; xia2021 . Recent measurements of neutron skin thickness from scattering of the spin-polarized electrons on 208Pb by the PREX-II experiments adhikari2021 and on 48C by the CREX adhikari2022 lead to controversy about the value of $L$ reed2021 ; Reinhard2021 ; reed2024 . Another experiment is the measurement of the spectra of charged pions produced by colliding rare isotope Sn beams with isotopically enriched Sn targets. The deduced value of $L$ has been shown to be 42 $<L<$ 117 MeV estee2021 . Also there are attempts to get information on $L$ from the astronomical observations associated with $X$-ray bursters sotani2018 . In the present work, the parameters, $V_{r}$, $c_{r}$, and $\lambda_{r}$ in $U_{\rm SJM2}$ [(43)] are fixed to be those used in SJM2 model in Ref. tnt2008 , so that the $E$ (TNR) is commonly given irrespective of choice of the other parameters. Allowable values of the $L$ are chosen so that the overall density-dependence of the energy contributions, $E$ (two-body), $E$ (TNA), and $E$ (total) around $\rho_{0}$ in SNM are similar to those obtained by the standard variational calculation by the use of the $v_{14}$ two-body potential with addition of the phenomenological TNR and TNA by Lagaris-Pandahripande [LP (1981)] lp1981 . It turns out that $E$ (two-body) in the present model is not bound for $L\lesssim 55$ MeV and that its density-dependence is quite different from that of LP (1981). For $L\gtrsim 75$ MeV, an onset density of KC exceeds a central density of neutron star with maximum mass. Hence the typical three values $L$ = (60, 65, 70) MeV are taken throughout this paper. In Table 2, the relevant quantities associated with the (MRMF+UTBR+TNA) model are listed for three cases of $L$ = (60, 65, 70) MeV.